# Sparse Alignment Enhanced Latent Diffusion Transformer for Zero-Shot Speech Synthesis

### v1

@@ 4.1 Experimental setup
4.2 Results of Zero-Shot Speech Synthesis
Evaluation Baselines.
We compare the zero-shot speech synthesis performance of S-DiT with 11 strong baselines, including: 1) VALL-E 2 (Chen et al., 2024a); 2) VoiceBox (Matthew et al., 2023); 3) DiTTo-TTS (Lee et al., 2024a); 4) NaturalSpeech 3 (Ju et al., 2024); 5) CosyVoice (Du et al., 2024); 6) MaskGCT (Wang et al., 2024); 7) F5-TTS (Chen et al., 2024b); 8) E2 TTS (Eskimez et al., 2024). Explanation and details of the selected baseline systems are provided in Appendix A.4.
Analysis
As shown in Table 1, we can see that 1) S-DiT achieves state-of-the-art SIM-O, SMOS, and WER scores, comparable to NaturalSpeech 3 (the counterpart with forced alignment), and significantly surpasses other baselines without explicit alignments. The improved SIM-O and SMOS suggest that the proposed sparse alignment effectively simplifies the text-to-speech mapping challenge like predefined forced duration information, allowing the model to focus more on learning timbre information. And the improved WER indicates that S-DiT also enjoys strong robustness; 2) S-DiT significantly surpasses all baselines in terms of CMOS, demonstrating the effectiveness of the proposed sparse alignment strategy; 3) After the PeRFlow acceleration, the student model of S-DiT shows on par quality with the teacher model and enjoys fast inference speed. We also conduct the experiments on the LibriSpeech-PC test-clean set provided by F5-TTS and the results are shown in Table 2, which also demonstrates that our method achieves state-of-the-art performance in terms of speaker similarity and speech intelligibility. The duration controllability of S-DiT is verified in Appendix E. In the demo page, we also demonstrate that our method can maintain high naturalness even when the performance of the duration predictor is suboptimal (while S-DiT with forced alignment fails).
GT - 45.1 0.591 0.783 4.03 +0.09 3.95
CTA-TTS 5.98 41.1 0.602 0.799 3.72 -0.60 3.64
S-DiT 5.69 42.3 0.601 0.790 3.84 +0.00 3.89
@@ 4.3 Experiments of Prosodic Naturalness
4.4 Results of Accented TTS
In this subsection, we evaluate the accented TTS performance of our model on the L2-ARCTIC dataset (Zhao et al., 2018). This corpus includes recordings from non-native speakers of English whose first languages are Hindi, Korean, etc. In this experiment, we focus on verifying whether our model and baseline can synthesize natural speech with different accent types (standard English or English with specific accents) while maintaining consistent vocal timbre. We compare our S-DiT model with CTA-TTS (Liu et al., 2024a). More details of the baseline model are provided in Appendix A.5. 1) First, we evaluate whether the models can synthesize high-quality speeches with accents. As shown in Table 4, our S-DiT model significantly outperforms the CTA-TTS baseline in terms of the subjective accent similarity MOS core, the MCD (dB) values, and the statistical moments ($\sigma$, $\gamma$, and $\kappa$) of pitch distributions. These results demonstrate the superior accent learning capability of S-DiT compared to the baseline system. Besides, the S-DiT model achieves higher CMOS and SMOS scores compared to CTA-TTS, indicating a significant improvement in speech quality and speaker similarity; 2) Secondly, we evaluate whether the models can accurately control the accent types of the generated speeches. We follow CTA-TTS to conduct the intensity classification experiment (Liu et al., 2024a). At run-time, we generate speeches with two accent types, and the listeners are instructed to classify the perceived accent categories, including “standard” and “accented”. Figure 2 shows that our S-DiT significantly surpasses CTA-TTS in terms of accent controllability.
@@ 4.4 Results of Accented TTS
Table 6: Comparison of zero-shot TTS performance of S-DiT using different speech compression models on the LibriSpeech test-clean set.
@@ 4.5 Evaluation of WaveVAE
Second, to demonstrate the impact of different speech compression models on the overall performance of the TTS system, we extracted the latents from Encodec and DAC, respectively, for training our S-DiT model. We report the experimental results in Table 6. It can be seen that our method outperforms “w/ DAC” and “w/ Encodec”, due to the fact that the latent space of our speech compression model is more compact (only 25 tokens per second). The results demonstrate the importance of our WaveVAE, a high-compression, high-reconstruction-quality speech codec model, for TTS systems. This conclusion is also verified by a previous work (Lee et al., 2024a), which shows compact target latents facilitate learning in diffusion models.
@@ 4.6 Ablation Studies
We test the following four settings: 1) w/ FA, which replaces the sparse alignment in S-DiT with forced alignment used in  (Matthew et al., 2023; Shen et al., 2023); 2) w/o A., we do not use the predefined alignments and modeling the duration information implicitly; 3) w/ CFG, we use the standard CFG following the common practice in Diffusion-based TTS; 4) w/o CFG, we do not use the CFG mechanism. All tests follow the experimental setup described in Section 4.2. The results are shown in Table 7. For settings 1) and 2), it can be observed that both forced alignment and sparse alignment can enhance the performance of speech synthesis models. However, compared to forced alignment, sparse alignment does not constrain the model’s search space, leading to a prosodic naturalness (see Section 4.3). Therefore, the sparse alignment strategy achieves $+0.17$ CMOS compared to the forced alignment strategy. For setting 3), compared with the standard CFG, our multi-condition CFG performs slightly better as it allows for flexible control over the weights between the text prompt and the speaker prompt. Setting 4) proves that the CFG mechanism is crucial for S-DiT. Additionally, we visualize the attention score matrices from different transformer layers in S-DiT in Appendix F, leading to some interesting observations.
@@ 5 Conclusions
In this paper, we introduce S-DiT, a zero-shot TTS framework that leverages novel sparse alignment boundaries to ease the difficulty of alignment learning while retaining the naturalness of the generated speeches. This strategy allows S-DiT to combine the strengths of methods with both implicit alignments and predefined hard alignments. Additionally, we employ the PeRFlow technique to further accelerate the generation process and design a multi-condition CFG strategy to offer more flexible control over accents. Experimental results show that S-DiT achieves state-of-the-art zero-shot TTS speech quality while maintaining a more efficient pipeline. Moreover, the sparse alignment strategy also shows enhanced prosodic naturalness and higher robustness against a suboptimal duration predictor. Due to space constraints, further discussions are provided in the appendix.
@@
  • Function Coverage. We can make S-DiT more user-friendly by enabling it to generate speech in various styles according to text descriptions through instruction-based fine-tuning. We can further fine-tune S-DiT on the paralinguistic corpus, allowing it to generate speech that is closer to a natural human style.
@@ A.1 Model Configuration
  • The S-DiT model use the standard transformer block from LLAMA (Dubey et al., 2024) as the basic structure, which comprises a 24-layer Transformer with 16 attention heads and 1024 embedding dimensions. It contains 339M parameters in total. We adopt the Rotary Position Embedding (RoPE) (Su et al., 2024) as the positional embedding following the common practice in LLAMA implementations. For simplicity, we do not use the phoneme encoder and style encoder like previous works. We only use a linear projection layer to transform these features to the same dimension. During training, we use 8 A100 80GB GPUs with a batch size of 10K latent frames per GPU for 1M steps. We use the Adam optimizer with a learning rate of $5\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warmup steps. In zero-shot TTS experiments, we set the text guidance scale $\alpha_{txt}$ and the speaker guidance scale $\alpha_{spk}$ to 2.5 and 3.5, respectively. In accented TTS experiments, we set $\alpha_{spk}=6.5$, $\alpha_{txt}=1.5$ to generate the accented speech and set $\alpha_{spk}=2.0$, $\alpha_{txt}=5.0$ to generate the speech with standard English.
@@ A.3 Sampling Strategy
For S-DiT, we applied the Euler sampler with a fixed step size following the common practice in flow ODE sampling. We use 25 and 8 sampling steps for S-DiT and S-DiT-accelerated, respectively.
@@ Appendix D Details about Data and Model Scaling Experiments
The data/model scalability is crucial for practical TTS systems. To evaluate the scalability of S-DiT in Section 4.6, we construct a 600kh internal multilingual training corpus comprising both English and Chinese speech. Most of the audiobook recordings are crawled from YouTube and online podcasts like novelfm [^5] https://novelfm.changdunovel.com/. We also include the academic datasets like LibriLight (Kahn et al., 2020), WenetSpeech (Zhang et al., 2022), and GigaSpeech (Chen et al., 2021). Since the crawled corpus may contain unlabelled speeches. We transcribe them using an internal ASR model.
@@
We scale up S-DiT from 0.5B to 7.0B following the hyper-parameter settings in Qwen 2 (Yang et al., 2024a). In this experiment, we only increase the parameters of the S-DiT model to verify its scalability. The parameters of the speech compression VAE remained unchanged. In theory, expanding the parameters of both models could yield the optimal results, which we leave for future work.
@@ Results
We evaluate the effectiveness of data and model scaling for the proposed S-DiT model. In this experiment, we train models with 0.5B parameters on multilingual internal datasets with data sizes of 2kh, 40kh, 200kh, and 600kh, respectively. We also train models with 0.5B, 1.5B, and 7.0B parameters on the 600kh dataset. We evaluate the zero-shot TTS performance in terms of speaker similarity (Sim-O) and speech intelligibility (WER) on an internal test set consisting of 400 speech samples from various sources. Based on Table 8, we conclude that: 1) as the data size increases from 2kh to 600kh, both the model’s speaker similarity and speech intelligibility improve consistently, demonstrating strong data scalability of our model; 2) as the model size scales from 0.5B to 7.0B parameters, SIM-O improves by 12.1% and WER decreases by 9.52%, validating the model scalability of S-DiT. Additionally, we find that increasing the model parameters enhances its para-linguistic capabilities, with specific audio examples available on the demo page.
@@
Appendix E Duration Controllability of S-DiT
In this section, we aim to verify S-DiT’s duration control capabilities through case studies. We randomly selected a speech prompt from the test set and used the sentence “Notably, raising questions about both the size of the perimeter and efforts to sweep and secure.” as the target sentence to generate speeches. In the generation process, we first control the sentence-level duration by multiplying the time coordinates of the phoneme anchors described in Section 3.2 by a fixed value. As shown in Figure 5, our S-DiT demonstrates good sentence-level duration control. Moreover, our S-DiT is also capable of fine-grained phoneme-level duration control. As illustrated in Figure 5, we multiplied the anchor coordinates of the phoneme within the red box by a fixed value while keeping the relative positions of other phoneme anchors unchanged. The figure shows that our S-DiT also exhibits good fine-grained phoneme-level duration controllability.
@@ Appendix E Duration Controllability of S-DiT
Figure 6: Visualization of Attention Matrices from different layers in S-DiT.
@@ Appendix F Visualization of Attention Matrices
We visualize the attention matrices from all layers in the 1.4B S-DiT model, using 8 sampling steps. From Figure 6, we observe: 1) within the same layer, despite different timesteps, the attention matrices remain identical. In other words, the function of each layer stays consistent across timesteps; 2) the functions of the transformer layers can be categorized into three types. As shown in Figure 6 (a), the bottom layers handle text and audio feature extraction; in Figure 6 (b), the middle layers focus on speech-text alignment; and in Figure 6 (c), the top layers refine the target latent features.
@@ Appendix H Experiments of Prosodic Naturalness for Zero-Shot TTS
We also conduct the ablation studies using the objective metrics MCD, SSIM, STOI, GPE, VDE, and FFE following InstructTTS (Yang et al., 2024c) to evaluate the prosodic naturalness of our proposed method. We conduct experiments on the LibriSpeech test-clean 2.2-hour subset (following the setup in VALL-E 2 and Voicebox). The results are shown in the Table below. We compare S-DiT with the following baselines: 1) “Ours w/ Forced Alignment”, we replace the sparse alignment with the forced alignment; 2) “Ours w/ Standard CFG”, we replace the multi-condition CFG with standard CFG; 3) “Ours w/ Standard AR Duration”, we replace the duration from F-LM with the duration from standard AR duration predictor following SimpleSpeech 2 (Yang et al., 2024b). The results in Table 9 show that sparse alignment brings significant improvements, and both multi-condition CFG and F-LM duration contribute positively to the performance.
@@ Appendix I Experiments with Longer Samples
To directly compare S-DiT’s robustness to long sequences against other AR models, we have conducted experiemnts for a test set with longer samples. Specifically, we randomly select 10 sentences, each containing more than 50 words. For each speaker in the LibriSpeech test-clean set, we randomly chose a 3-second clip as a prompt, resulting in 400 target samples in total. To make our results more convincing, we include strong-performing TTS models, VoiceCraft (Peng et al., 2024) and CosyVoice (AR+NAR) (Du et al., 2024), as our baselines. The results for longer samples are presented in Table 10. As shown, compared to the baseline systems, S-DiT does not exhibit a significant decline in speech intelligibility when generating longer sentences, illustrating the effectiveness of the combination of F-LM and S-DiT.
@@
The proposed model, S-DiT, is designed to advance zero-shot TTS technologies, making it easier for users to generate personalized speech. When used responsibly and legally, this technique can enhance applications such as movies, games, podcasts, and various other services, contributing to increasing convenience in everyday life. However, we acknowledge the potential risks of misuse, such as voice cloning for malicious purposes. To mitigate this risk, solutions like building a corresponding deepfake detection model will be considered. Additionally, we plan to incorporate watermarks and verification methods for synthetic audio to ensure ethical use in real-world applications. Restrictions will also be included in the licensing of our project to further prevent misuse. By addressing these ethical concerns, we aim to contribute to the development of responsible and beneficial AI technologies, while remaining conscious of the potential risks and societal impact.
@@
We have taken several steps to ensure the reproducibility of the experiments and results presented in this paper: 1) the architecture and algorithm of the S-DiT model are described in Section 3 and and relevant hyperparameters are fully described in Appendix A.1; 2) The evaluation metrics, including WER, SIM-O, MCD (dB), the moments of the pitch distribution, alignment error, CMOS, SMOS, and ASMOS, are described in detail in Section 4.1; 3) For most of the key experiments, we utilize publicly available datasets such as LibriLight, LibriSpeech, and L2Arctic. The selection of the test sets is identical to that used in previous zero-shot TTS research. However, as the publicly available datasets are insufficient for our data scaling experiments, we construct a larger dataset, which is described in detail in Appendix D; 4) To ensure reproducibility of the results, we have carefully set random seeds in our experiments and the random seeds are provided in Appendix A.2. All objective results reported are based on the average performance across multiple runs.

### v4

@@ 4.3 Experiments of Prosodic Naturalness
4.4 Results of Accented TTS
In this subsection, we evaluate the accented TTS performance of our model on the L2-ARCTIC dataset (Zhao et al., 2018). This corpus includes recordings from non-native speakers of English whose first languages are Hindi, Korean, etc. In this experiment, we focus on verifying whether our model and baseline can synthesize natural speech with different accent types (standard English or English with specific accents) while maintaining consistent vocal timbre. We compare our MegaTTS 3 model with CTA-TTS (Liu et al., 2024a). More details of the baseline model are provided in Appendix A.5. 1) First, we evaluate whether the models can synthesize high-quality speeches with accents. As shown in Table 3, our MegaTTS 3 model significantly outperforms the CTA-TTS baseline in terms of the subjective accent similarity MOS core, the MCD (dB) values, and the statistical moments ($\sigma$, $\gamma$, and $\kappa$) of pitch distributions. These results demonstrate the superior accent learning capability of MegaTTS 3 compared to the baseline system. Besides, the MegaTTS 3 model achieves higher CMOS and SMOS scores compared to CTA-TTS, indicating a significant improvement in speech quality and speaker similarity; 2) Secondly, we evaluate whether the models can accurately control the accent types of the generated speeches. We follow CTA-TTS to conduct the intensity classification experiment (Liu et al., 2024a). At run-time, we generate speeches with two accent types, and the listeners are instructed to classify the perceived accent categories, including “standard” and “accented”. Figure 2 shows that our MegaTTS 3 significantly surpasses CTA-TTS in terms of accent controllability.
@@ 4.4 Results of Accented TTS
Table 6: Comparison of zero-shot TTS performance of MegaTTS 3 using different speech compression models on the LibriSpeech test-clean set.
@@ 4.5 Evaluation of WaveVAE
Second, to demonstrate the impact of different speech compression models on the overall performance of the TTS system, we extracted the latents from Encodec and DAC, respectively, for training our MegaTTS 3 model. We report the experimental results in Table 6. It can be seen that our method outperforms “w/ DAC” and “w/ Encodec”, due to the fact that the latent space of our speech compression model is more compact (only 25 tokens per second). The results demonstrate the importance of our WaveVAE, a high-compression, high-reconstruction-quality speech codec model, for TTS systems. This conclusion is also verified by a previous work (Lee et al., 2024a), which shows compact target latents facilitate learning in diffusion models.
@@ 4.6 Ablation Studies
We test the following four settings: 1) w/ FA, which replaces the sparse alignment in MegaTTS 3 with forced alignment used in  (Matthew et al., 2023; Shen et al., 2023); 2) w/o A., we do not use the predefined alignments and modeling the duration information implicitly; 3) w/ CFG, we use the standard CFG following the common practice in Diffusion-based TTS; 4) w/o CFG, we do not use the CFG mechanism. All tests follow the experimental setup described in Section 4.2. The results are shown in Table 7. For settings 1) and 2), it can be observed that both forced alignment and sparse alignment can enhance the performance of speech synthesis models. However, compared to forced alignment, sparse alignment does not constrain the model’s search space, leading to a prosodic naturalness (see Section 4.3). Therefore, the sparse alignment strategy achieves $+0.17$ CMOS compared to the forced alignment strategy. For setting 3), compared with the standard CFG, our multi-condition CFG performs slightly better as it allows for flexible control over the weights between the text prompt and the speaker prompt. Setting 4) proves that the CFG mechanism is crucial for MegaTTS 3. Additionally, we visualize the attention score matrices from different transformer layers in MegaTTS 3 in Appendix F, leading to some interesting observations.
@@ 5 Conclusions
In this paper, we introduce MegaTTS 3, a zero-shot TTS framework that leverages novel sparse alignment boundaries to ease the difficulty of alignment learning while retaining the naturalness of the generated speeches. This strategy allows MegaTTS 3 to combine the strengths of methods with both implicit alignments and predefined hard alignments. Additionally, we employ the PeRFlow technique to further accelerate the generation process and design a multi-condition CFG strategy to offer more flexible control over accents. Experimental results show that MegaTTS 3 achieves state-of-the-art zero-shot TTS speech quality while maintaining a more efficient pipeline. Moreover, the sparse alignment strategy also shows enhanced prosodic naturalness and higher robustness against a suboptimal duration predictor. Due to space constraints, further discussions are provided in the appendix.
@@
  • Function Coverage. We can make MegaTTS 3 more user-friendly by enabling it to generate speech in various styles according to text descriptions through instruction-based fine-tuning. We can further fine-tune MegaTTS 3 on the paralinguistic corpus, allowing it to generate speech that is closer to a natural human style.
@@ A.1 Model Configuration
  • The MegaTTS 3 model use the standard transformer block from LLAMA (Dubey et al., 2024) as the basic structure, which comprises a 24-layer Transformer with 16 attention heads and 1024 embedding dimensions. It contains 339M parameters in total. We adopt the Rotary Position Embedding (RoPE) (Su et al., 2024) as the positional embedding following the common practice in LLAMA implementations. For simplicity, we do not use the phoneme encoder and style encoder like previous works. We only use a linear projection layer to transform these features to the same dimension. During training, we use 8 A100 80GB GPUs with a batch size of 10K latent frames per GPU for 1M steps. We use the Adam optimizer with a learning rate of $5\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warmup steps. In zero-shot TTS experiments, we set the text guidance scale $\alpha_{txt}$ and the speaker guidance scale $\alpha_{spk}$ to 2.5 and 3.5, respectively. In accented TTS experiments, we set $\alpha_{spk}=6.5$, $\alpha_{txt}=1.5$ to generate the accented speech and set $\alpha_{spk}=2.0$, $\alpha_{txt}=5.0$ to generate the speech with standard English.
@@ A.3 Sampling Strategy
For MegaTTS 3, we applied the Euler sampler with a fixed step size following the common practice in flow ODE sampling. We use 25 and 8 sampling steps for MegaTTS 3 and MegaTTS 3-accelerated, respectively.
@@ Appendix D Details about Data and Model Scaling Experiments
The data/model scalability is crucial for practical TTS systems. To evaluate the scalability of MegaTTS 3 in Section 4.6, we construct a 600kh internal multilingual training corpus comprising both English and Chinese speech. Most of the audiobook recordings are crawled from YouTube and online podcasts like novelfm [^5] https://novelfm.changdunovel.com/. We also include the academic datasets like LibriLight (Kahn et al., 2020), WenetSpeech (Zhang et al., 2022), and GigaSpeech (Chen et al., 2021). Since the crawled corpus may contain unlabelled speeches. We transcribe them using an internal ASR model.
@@
We scale up MegaTTS 3 from 0.5B to 7.0B following the hyper-parameter settings in Qwen 2 (Yang et al., 2024a). In this experiment, we only increase the parameters of the MegaTTS 3 model to verify its scalability. The parameters of the speech compression VAE remained unchanged. In theory, expanding the parameters of both models could yield the optimal results, which we leave for future work.
@@ Results
We evaluate the effectiveness of data and model scaling for the proposed MegaTTS 3 model. In this experiment, we train models with 0.5B parameters on multilingual internal datasets with data sizes of 2kh, 40kh, 200kh, and 600kh, respectively. We also train models with 0.5B, 1.5B, and 7.0B parameters on the 600kh dataset. We evaluate the zero-shot TTS performance in terms of speaker similarity (Sim-O) and speech intelligibility (WER) on an internal test set consisting of 400 speech samples from various sources. Based on Table 8, we conclude that: 1) as the data size increases from 2kh to 600kh, both the model’s speaker similarity and speech intelligibility improve consistently, demonstrating strong data scalability of our model; 2) as the model size scales from 0.5B to 7.0B parameters, SIM-O improves by 12.1% and WER decreases by 9.52%, validating the model scalability of MegaTTS 3. Additionally, we find that increasing the model parameters enhances its para-linguistic capabilities, with specific audio examples available on the demo page.
@@
Appendix E Duration Controllability of MegaTTS 3
In this section, we aim to verify MegaTTS 3’s duration control capabilities through case studies. We randomly selected a speech prompt from the test set and used the sentence “Notably, raising questions about both the size of the perimeter and efforts to sweep and secure.” as the target sentence to generate speeches. In the generation process, we first control the sentence-level duration by multiplying the time coordinates of the phoneme anchors described in Section 3.2 by a fixed value. As shown in Figure 5, our MegaTTS 3 demonstrates good sentence-level duration control. Moreover, our MegaTTS 3 is also capable of fine-grained phoneme-level duration control. As illustrated in Figure 5, we multiplied the anchor coordinates of the phoneme within the red box by a fixed value while keeping the relative positions of other phoneme anchors unchanged. The figure shows that our MegaTTS 3 also exhibits good fine-grained phoneme-level duration controllability.
@@ Appendix E Duration Controllability of MegaTTS 3
Figure 6: Visualization of Attention Matrices from different layers in MegaTTS 3.
@@ Appendix F Visualization of Attention Matrices
We visualize the attention matrices from all layers in the 1.4B MegaTTS 3 model, using 8 sampling steps. From Figure 6, we observe: 1) within the same layer, despite different timesteps, the attention matrices remain identical. In other words, the function of each layer stays consistent across timesteps; 2) the functions of the transformer layers can be categorized into three types. As shown in Figure 6 (a), the bottom layers handle text and audio feature extraction; in Figure 6 (b), the middle layers focus on speech-text alignment; and in Figure 6 (c), the top layers refine the target latent features.
@@ Appendix H Experiments of Prosodic Naturalness for Zero-Shot TTS
We also conduct the ablation studies using the objective metrics MCD, SSIM, STOI, GPE, VDE, and FFE following InstructTTS (Yang et al., 2024c) to evaluate the prosodic naturalness of our proposed method. We conduct experiments on the LibriSpeech test-clean 2.2-hour subset (following the setup in VALL-E 2 and Voicebox). The results are shown in the Table below. We compare MegaTTS 3 with the following baselines: 1) “Ours w/ Forced Alignment”, we replace the sparse alignment with the forced alignment; 2) “Ours w/ Standard CFG”, we replace the multi-condition CFG with standard CFG; 3) “Ours w/ Standard AR Duration”, we replace the duration from F-LM with the duration from standard AR duration predictor following SimpleSpeech 2 (Yang et al., 2024b). The results in Table 9 show that sparse alignment brings significant improvements, and both multi-condition CFG and F-LM duration contribute positively to the performance.
@@ Appendix I Experiments with Longer Samples
To directly compare MegaTTS 3’s robustness to long sequences against other AR models, we have conducted experiemnts for a test set with longer samples. Specifically, we randomly select 10 sentences, each containing more than 50 words. For each speaker in the LibriSpeech test-clean set, we randomly chose a 3-second clip as a prompt, resulting in 400 target samples in total. To make our results more convincing, we include strong-performing TTS models, VoiceCraft (Peng et al., 2024) and CosyVoice (AR+NAR) (Du et al., 2024), as our baselines. The results for longer samples are presented in Table 10. As shown, compared to the baseline systems, MegaTTS 3 does not exhibit a significant decline in speech intelligibility when generating longer sentences, illustrating the effectiveness of the combination of F-LM and MegaTTS 3.
@@
The proposed model, MegaTTS 3, is designed to advance zero-shot TTS technologies, making it easier for users to generate personalized speech. When used responsibly and legally, this technique can enhance applications such as movies, games, podcasts, and various other services, contributing to increasing convenience in everyday life. However, we acknowledge the potential risks of misuse, such as voice cloning for malicious purposes. To mitigate this risk, solutions like building a corresponding deepfake detection model will be considered. Additionally, we plan to incorporate watermarks and verification methods for synthetic audio to ensure ethical use in real-world applications. Restrictions will also be included in the licensing of our project to further prevent misuse. By addressing these ethical concerns, we aim to contribute to the development of responsible and beneficial AI technologies, while remaining conscious of the potential risks and societal impact.
@@
We have taken several steps to ensure the reproducibility of the experiments and results presented in this paper: 1) the architecture and algorithm of the MegaTTS 3 model are described in Section 3 and and relevant hyperparameters are fully described in Appendix A.1; 2) The evaluation metrics, including WER, SIM-O, MCD (dB), the moments of the pitch distribution, alignment error, CMOS, SMOS, and ASMOS, are described in detail in Section 4.1; 3) For most of the key experiments, we utilize publicly available datasets such as LibriLight, LibriSpeech, and L2Arctic. The selection of the test sets is identical to that used in previous zero-shot TTS research. However, as the publicly available datasets are insufficient for our data scaling experiments, we construct a larger dataset, which is described in detail in Appendix D; 4) To ensure reproducibility of the results, we have carefully set random seeds in our experiments and the random seeds are provided in Appendix A.2. All objective results reported are based on the average performance across multiple runs.